Can you please help me solving this question. Kaye sent an invitation to Lara saying "I'm inviting you to my birthday party! My age will be 6 years less than thrice my age ten years ago. How old will I be on my birthday.

Answer by ankor@dixie-net.com(22740)

let a = K's present age
:
" My age will be 6 years less than thrice my age ten years ago."
a = 3(a-10) - 6
a = 3a - 30 - 6
a = 3a - 36
36 = 3a - a
36 = 2a
a = 36/2
a = 18 yrs
:
Check, 10 yrs ago she was 8, 3(8) - 6 = 18